OpenStudy (anonymous):

Solve. (x + 4)^2 = 4 x = 0 and x = –4 x = –2 and x = –6 x = 20 and x = 12 x = 0 and x = 4

OpenStudy (anonymous):

x=-2 , -6

OpenStudy (anonymous):

(x+4)^2=4 =>x^2+8x+16=4 [since, (a+b)^2=a^2+b^2+2ab, by formula] =>x^2+8x+16-4=0 By factorising =>x^2+8x+12=0 =>(x+2)(x+6)=0 =>x+2=0|x+6=0 =>x=-2|x=-6 Ans:x=-2,-6